# Contrastive explainable clustering with differential privacy

### v1

@@ Abstract
This paper presents a novel approach in Explainable AI (XAI), integrating contrastive explanations with differential privacy in clustering methods. For several basic clustering problems, including $k$-median and $k$-means, we give efficient differential private contrastive explanations that achieve essentially the same explanations as those that non-private clustering explanations can obtain. We define contrastive explanations as the utility difference between the original clustering utility and utility from clustering with a specifically fixed centroid. In each contrastive scenario, we designate a specific data point as the fixed centroid position, enabling us to measure the impact of this constraint on clustering utility under differential privacy. Extensive experiments across various datasets show our method’s effectiveness in providing meaningful explanations without significantly compromising data privacy or clustering utility. This underscores our contribution to privacy-aware machine learning, demonstrating the feasibility of achieving a balance between privacy and utility in the explanation of clustering tasks.
@@ 1 Introduction
Different notions of clustering are fundamental primitives in several areas, including machine learning, data science, and operations research Reddy (2018). Clustering objectives (e.g., $k$-median and $k$-means), crucial in facility location planning, have been notably applied in pandemic response strategies, such as determining optimal sites for COVID-19 vaccination and testing centers  Bertsimas et al. (2022); Mehrab et al. (2022). The distance to a cluster center becomes a significant factor in these scenarios, particularly highlighted during the COVID-19 pandemic, where accessible vaccine locations played a pivotal role  Rader et al. (2022). This context often leads to individuals, especially those distant from a cluster center, seeking explanations for their specific assignment or location in the clustering plan. For example, consider a resident questioning why a COVID-19 testing center isn’t located next to their home. Such queries reflect the need for explainability in clustering decisions, where factors like population density, resource allocation, and public health considerations impact site selection. The resident’s inquiry underscores the importance of transparent methodologies in clustering that can provide understandable explanations for facility placements, especially in critical health scenarios.
Such questions fall within the area of Explainable AI, which is a rapidly growing and vast area of research Finkelstein et al. (2022); Sreedharan et al. (2020); Miller (2019); Boggess et al. (2023). We focus on post-hoc explanations, especially contrastive explanations, e.g., Miller (2019), which address “why P instead of Q?” questions. Such methods have been used in many applications, including multi-agent systems, reinforcement learning, and contrastive analysis, e.g., Boggess et al. (2023); Sreedharan et al. (2021); Madumal et al. (2020); Sreedharan et al. (2020). In the context of the clustering problems mentioned above, a possible explanation for a resident is the increase in the cost of the clustering if a center is placed close to the resident, compared to the cost of an optimal solution, which is relatively easy to compute. This increased cost reflects a trade-off, where optimizing for one resident’s convenience could result in worse centroid positions for other residents, highlighting the complex decision-making process in clustering for fair resource distribution.
@@
Data privacy is a crucial concern across various fields, and Differential Privacy (DP) is one of the most widely used and rigorous models for privacy Dwork and Roth (2014). We focus on the setting where the set of data points $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ are private; for instance, in the vaccine center deployment problem, each $x_{i}$ corresponds to an individual who is seeking vaccines, and would like to be private. There has been a lot of work on the design of differentially private solutions to clustering problems like $k$-median and $k$-means under such a privacy model Ghazi et al. (2020); Gupta et al. (2010); Balcan et al. (2017); Stemmer and Kaplan (2018). These methods guarantee privacy of the data points $\mathbf{X}$ and output a set of centers $c_{1},\ldots,c_{k}$ (which are public) such that its cost, denoted by $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ is within multiplicative and additive factors $w$ and $t$ of the optimal solution $OPT_{\mathbf{X}}$, i.e., $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT_{\mathbf{X}}+t$.
@@
While there has been significant progress in various domains of differential privacy, the intersection of explainability and differential privacy still needs to be explored. In clustering problems, a potential private contrastive explanation given to agent $i$ is to provide the $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, where $S_{\epsilon}$ is a set of centers generated by private algorithm (with privacy budget $\epsilon$), and $\text{cost}(S^{(i)}_{\epsilon})$ is the cost of the clustering solution while fixing one center at agent $i$’s location. However, giving such a private contrastive explanation to each agent $i$ naively using a $(w,t)$-approximate private clustering would lead to $(w,t\sqrt{n}/\epsilon)$-approximation, even if we use advanced composition techniques for privacy. The central question of our research: is it possible to offer each user a private contrastive explanation with a limited overall privacy budget?
@@
2. We present an $\epsilon$-DP mechanism, PrivateExplanations, which provides a contrastive explanation to each agent while ensuring the same utility bounds as private clustering in Euclidean spaces Ghazi et al. (2020). We use the private coreset technique of Ghazi et al. (2020), which is an intermediate private data structure that preserves similar clustering costs as the original data. The main technical challenge is to derive rigorous bounds on the approximation factors for all the contrastive explanations.
3. We evaluate our methods empirically on realistic datasets used in the problems of vaccine center deployment in Virginia Mehrab et al. (2022). Our results show that the privacy-utility trade-offs are similar to those in private clustering, and the errors are quite low even for reasonable privacy budgets, demonstrating the effectiveness of our methods.
Our research stands out by seamlessly integrating differential privacy into contrastive explanations, maintaining the quality of explanations even under privacy constraints. This work bridges the gap between privacy and explainability, marking a significant advancement in privacy-aware machine learning. Due to space limitations, we present many technical details and experimental evaluations in the Appendix.
@@ 2 Related Work
Our work considers differential privacy for explainable AI in general (XAI) and Multi-agent explanations (XMASE) in particular, focusing on post-hoc contrastive explanations for clustering. We summarize some of the work directly related to our paper; additional discussion is presented in the Appendix, due to space limitations.
Extensive experiments presented in Saifullah et al. (2022) demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy.
@@ 3 Preliminaries
Let $\mathbf{X}\subset\mathbb{R}^{d}$ denote a dataset consisting of $d$-dimensional points (referred to sometimes as agents). We consider the notion of $(k,p)$-clustering, as defined by Ghazi et al. (2020).
@@
($(w,t)$-approximation). Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, let $OPT^{p,k}_{\mathbf{X}}=\min_{c_{1},\ldots,c_{k}\in\mathbb{R}^{d}}\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ denote the cost of an optimal $(k,p)$-clustering. We say $c_{1},\ldots,c_{k}$ is a $(w,t)$– approximation to a $(k,p)$-optimal clustering if $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT^{p,k}_{\mathbf{X}}+t$.
@@
We operate under the assumption that the data points in $\mathbf{X}$ represent the private information of users or agents. We say $\mathbf{X},\mathbf{X}^{\prime}$ are neighboring (denoted by $\mathbf{X}\sim\mathbf{X}^{\prime}$) if they differ in one data point. When a value is disclosed to an individual agent $i$, it is imperative to treat the remaining clients in $\mathbf{X}-\{i\}$ as private entities. To address this necessity, we adopt the following privacy model:
@@
A mechanism $\mathcal{M}$ is $\epsilon$-$i$-exclusive DP if, $\forall X,X^{\prime}:X\setminus\{i\}\sim X^{\prime}\setminus\{i\}$, and for all $S\subseteq Range(\mathcal{M})$:
@@
Note that if $\mathcal{M}$ is $\epsilon$-DP, it is $\epsilon$-$i$-exclusive DP for any $i$.
Relation to Joint-DP model. Kearns et al. (2014) defines Joint-DP–a privacy model that “for each of agent $i$, knowledge of the $n-1$ agents does not reveal much about agent $i$”. It contrasts with our privacy model, where each explanation to agent $i$ uses direct and private knowledge of agent $i$ (that private knowledge will be returned to only agent $i$) and such knowledge does not reveal much about $n-1$ other agents.
In clustering scenarios such as k-means or k-median, it is beneficial for an individual data point to be positioned near the cluster center, as this reduces its distance. This concept has practical applications in fields such as urban planning and public health. For example, clustering algorithms like $k$-means and $k$-median have been utilized to optimize the placement of mobile vaccine centers, as noted in  Mehrab et al. (2022). However, once the cluster centers are calculated, residents might wonder why a vaccine clinic has not been positioned nearer to their homes. Our objective is to provide explanations for each data point that is not near a cluster center. Following the approach suggested by  Miller (2019), one way to provide an explanation is to re-run the clustering algorithm, this time setting the resident’s location as a fixed centroid. This allows us to explain the changes in clustering costs, where an increase could describe why certain placements are less optimal.
The focus of this paper is twofold: residents not only need explanations but also wish to maintain their privacy, preferring not to disclose their precise locations/data publicly. Traditional private clustering methods struggle to offer these private explanations because they accumulate the privacy budget by continuously applying private clustering to answer numerous inquiries, ultimately resulting in explanations that are not private. Thus, we have developed a framework that delivers the necessary explanations to residents while protecting their confidentiality.
In essence, our framework is data-agnostic and capable of offering contrastive explanations for any private data within clustering algorithms, effectively balancing the need for privacy with the demand for transparency in decision-making.
@@
Private and Explainable Clustering problem (PrivEC) Given an instance $\mathbf{X}\in\mathcal{X}^{n}$, clustering parameters $k,p$, and a subset of data points: $V_{s}\subset\mathbf{X}$, the goal is to output:
@@
Explainable: For each $i\in V_{s}$, output $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$, which is $\epsilon$-$i$-exclusive DP.
$S^{(i)}_{\epsilon}$ is a private solution computed by the clustering algorithm while fixing one centroid to the position of agent $i$.
We assume that $S^{(i)}_{\epsilon}$ is not revealed to any agent, but $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$ is released to agent $i$ as contrastive explanation.

### v2

@@ Abstract
This paper presents a novel approach to Explainable AI (XAI) that combines contrastive explanations with differential privacy for clustering algorithms. Focusing on k-median and k-means problems, we calculate contrastive explanations as the utility difference between original clustering and clustering with a centroid fixed to a specific data point. This method provides personalized insights into centroid placement. Our key contribution is demonstrating that these differentially private explanations achieve essentially the same utility bounds as non-private explanations. Experiments across various datasets show that our approach offers meaningful, privacy-preserving, and individually relevant explanations without significantly compromising clustering utility. This work advances privacy-aware machine learning by balancing data protection, explanation quality, and personalization in clustering tasks.
@@ 1 Introduction
Different notions of clustering are fundamental primitives in several areas, including machine learning, data science, and operations research Reddy (2018). $k$-means and $k$-median clustering remain among the most important and widely used approaches, as demonstrated by recent advances in explainability, privacy, fairness, and contrastive learning Ghadiri et al. (2021); Moshkovitz et al. (2020); Zehtabi et al. (2024); Newling and Fleuret (2016). These problems often involve significant trade-offs between accessibility, resource allocation, and overall cost. For example, in emergency response planning, authorities must decide the optimal locations for ambulance stations to minimize response times across a city. Residents might question why an ambulance station isn’t located closer to their neighborhood, especially if they feel imbalances in resource distribution. Similarly, in retail, customers might wonder why a new store is not placed near their area, despite being part of a high-demand demographic. Explainability in these contexts provides transparency into how and why certain decisions are made, addressing questions like: ”Why was this location chosen instead of another?” This is particularly important in applications where the consequences of clustering decisions directly affect individuals or communities. Bobek et al. (2022); Ofek and Somech (2024)
Such questions fall within the area of Explainable AI, which is a rapidly growing and vast area of research Finkelstein et al. (2022); Sreedharan et al. (2020); Miller (2019); Boggess et al. (2023); Schleibaum et al. (2024). We focus on post-hoc explanations, especially contrastive explanations, e.g., Miller (2019); Pozanco et al. (2022), which address “why P instead of Q?” questions. For example, in warehouse optimization, contrastive explanations clarify why a specific location was chosen as a distribution center, considering constraints like storage capacity or demand Zehtabi et al. (2024). These methods are widely applied in multi-agent systems, reinforcement learning, and contrastive analysis Boggess et al. (2023, 2022); Sreedharan et al. (2021); Madumal et al. (2020); Sreedharan et al. (2020). In reinforcement learning, they explain actions by highlighting trade-offs, such as long-term rewards or risks van der Waa et al. (2018).
Following the approach introduced in Zehtabi et al. (2024); Pozanco et al. (2022), we explain clustering decisions by comparing the costs in two scenarios: $cost(S)$, the cost of an optimal clustering solution on the whole dataset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ and $cost(S^{(i)})$, the cost of a modified solution where we fix a center at a desired location requested by agent $x_{i}\in\mathbf{X}$. We explain the decision by showing $cost(S^{(i)})-cost(S)$, i.e., how much the overall clustering cost increases when we force a center to be in a specific location. A higher clustering cost indicates worse performance. This comparison reveals the trade-offs in clustering: optimizing for one specific location often leads to a higher overall cost, meaning a worse solution for everyone else. By examining the difference between the optimal clustering cost and the cost of the forced fixed clustering, we can understand why the algorithm chose certain locations for centers and not others. This approach helps people grasp the complex balancing act involved in clustering decisions, especially when trying to distribute resources fairly Zehtabi et al. (2024).
@@
Data privacy is a crucial concern across various fields, and Differential Privacy (DP) is one of the most widely used and rigorous models for privacy Dwork and Roth (2014). We focus on the setting where the set of data points $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ are private; for instance, in the ambulance center deployment problem, each $x_{i}$ represents an individual requiring emergency services and seeking to keep their information private. There has been a lot of work on the design of differentially private solutions to clustering problems such as $k$-median and $k$-means in such a privacy model Ghazi et al. (2020); Gupta et al. (2010); Balcan et al. (2017); Stemmer and Kaplan (2018).
@@
While there has been significant progress in various domains of differential privacy, the intersection of explainability and differential privacy still needs to be explored. In clustering problems, building on the formalization of explanations for combinatorial problems we provide a private contrastive explanation to agent $x_{i}$ by computing $\text{cost}(S^{(i)}_{\epsilon})-\text{cost}(S_{\epsilon})$. Here, $S_{\epsilon}$ represents a private solution using the privacy budget $\epsilon$, while $S^{(i)}_{\epsilon}$ denotes a private solution, when we constrain one center to be at a location specified by agent $x_{i}$ (this can be any point of interest, not necessarily the agent’s own location). This difference quantifies how the clustering cost changes when accommodating agent $i$’s position, offering a privacy-preserving explanation of the clustering decision. However, giving such a private contrastive explanation to each agent $i$ naively using a private clustering algorithm would require a high privacy budget due to composition, which impact the accuracy, and lead to misleading or uninformative results. The central question of our research: is it possible to offer each user an informative private contrastive explanation with a limited overall privacy budget?
@@
2. We present an $\epsilon$-DP mechanism, PrivateExplanations, which provides a contrastive explanation to each agent while ensuring the same utility bounds as private clustering in Euclidean spaces, offering personalized insights without compromising privacy or clustering quality. We use the private coreset technique of Ghazi et al. (2020), which is an intermediate private data structure that preserves similar clustering costs as the original data.
3. We evaluate our methods on diverse datasets with varying distributions and feature dimensions. Our results demonstrate privacy-utility trade-offs comparable to private clustering, with low clustering errors even at reasonable privacy budgets, showcasing the effectiveness of our approach. Our research stands out by seamlessly integrating differential privacy into contrastive explanations, maintaining the quality of explanations even under privacy constraints. This work bridges the gap between privacy and explainability, marking a significant advancement in privacy-aware machine learning. A key technical contribution of our work is the derivation of rigorous bounds on the approximation factors for all contrastive explanations, ensuring their reliability and effectiveness. Due to space limitations, we only present major technical details in the main paper. We maintain a full, updated version of this paper with complete proofs and extended experimental results at Nguyen et al. (2024).
@@ 2 Related Work
Our work considers differential privacy for explainable AI in general (XAI) and Multi-agent explanations (XMASE) in particular, focusing on post-hoc contrastive explanations for clustering. We summarize some of the work directly related to our paper; additional discussion is presented in the Appendix, due to space limitations. Extensive experiments presented in Saifullah et al. (2022) demonstrate non-negligible changes in explanations of black-box ML models through the introduction of privacy.
@@ 3 Preliminaries
Let $\mathbf{X}\subset\mathbb{R}^{d}$ denote a dataset consisting of $d$-dimensional points (referred as agents). We consider the notion of $(k,p)$-clustering, as defined by Definition 1.
@@
($(w,t)$-approximation). Given $k\in\mathbb{N}$, and a multiset $\mathbf{X}=\{x_{1},\ldots,x_{n}\}$ of points in the unit ball, let $OPT^{p,k}_{\mathbf{X}}=\min_{c_{1},\ldots,c_{k}\in\mathbb{R}^{d}}\\ \text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})$ denote the cost of an optimal $(k,p)$-clustering. We say $c_{1},\ldots,c_{k}$ is a $(w,t)$– approximation to a $(k,p)$-optimal clustering if $\text{cost}^{p}_{\mathbf{X}}(c_{1},\ldots,c_{k})\leq w\cdot OPT^{p,k}_{\mathbf{X}}+t$.
Let $OPT$ denote the cost of the optimal $(k,p)$-clustering, $OPT_{i}$ denote the cost of the optimal $(k,p)$-clustering, with a center fixed at position $z_{i}$ (the location chosen by agent $i$) and the remaining $k$-1 centers are chosen to optimize the objective. Let $w^{\prime},w^{\prime\prime}$ denote the maximum approximation (w.r.t. $OPT$ and $OPT_{i}$ respectively) of non-private clustering algorithms. These factors will be specified in Sections 4.1 and 4.2.
@@
We assume that the data points in $\mathbf{X}$ (i.e., users) are private, and say $\mathbf{X},\mathbf{X}^{\prime}$ are neighboring (denoted by $\mathbf{X}\sim\mathbf{X}^{\prime}$) if they differ in one data point. When a value is disclosed to an individual agent $i$, it is imperative to treat the remaining clients in $\mathbf{X}-\{i\}$ as private entities.
@@
A mechanism $\mathcal{M}$ is $\epsilon$-$i$-exclusion DP if, $\forall X,X^{\prime}:i\in X,i\in X^{\prime},X\setminus\{i\}\sim X^{\prime}\setminus\{i\}$, and for all $S\subseteq Range(\mathcal{M})$:
@@
We extend this to say that $\mathcal{M}$ is $\epsilon$-$Y$-exclusion DP if the above holds $\forall X,X^{\prime}:Y\subset X,Y\subset X^{\prime},X\setminus Y\sim X^{\prime}\setminus Y$.
We now define the PrivEC problem for providing private contrastive explanations, where each agent $x_{i}$ seeks an explanation for a center fixed at a location of their choosing denoted by $z_{i}$.
@@
Private and Explainable Clustering problem (PrivEC) Given an instance $\mathbf{X}\subset\mathbb{R}^{d}$, clustering parameters $k,p$, and a contrastive set of points $Z\subset\mathbb{R}^{d}$, the goal is to output:
@@
Explainable: For each agent $x_{i}\in X$, output $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$.
$S^{(i)}_{\epsilon}$ is a private solution computed by the clustering algorithm with one centroid fixed at the position requested by agent $i$.
We assume that $S^{(i)}_{\epsilon}$ is not revealed to any agent, but $\text{cost}(S^{(i)}_{\epsilon})$ $\text{cost}(S_{\epsilon})$ is released to agent $i$ as contrastive explanation, which is $\epsilon$-$i$-exclusion DP.
Lemma 1.
With probability at least $1-\beta$, cost($S_{\epsilon}$) (clustering cost) is a $(w,t)$-approximation of $OPT$, where [^1] We use the notation $\mathcal{O}_{p,\alpha}$ to explicitly ignore factors of $p,\alpha$:
$w$ $=w^{\prime}(1+\alpha),$
$t$ $=w^{\prime}\mathcal{O}_{p,\alpha}\left((k/\beta)^{O_{p,\alpha}(1)}.\text{polylog}(n/\beta)/\epsilon\right)$
$\alpha$ is the approximation parameter for the utility of clustering and explanations. $\beta$ is the failure probability of the utility guarantees of clustering and explanations.
Lemma 2.
DP of fixed centroid yield additional cost. Fix an $i$. If $OPT_{i}\geq w^{\prime\prime}(1+\alpha)OPT+t^{(i)}$, then with probability at least $1-2\beta$, cost($S_{\epsilon}$) and cost($S^{(i)}_{\epsilon}$) computed by Algorithm peCluster satisfies that $cost(S^{(i)}_{\epsilon})>cost(S_{\epsilon}$).